Carmen is 14 years older than Sarah.
Five years ago, Carmen was three times as old is Sarah was then.

Answer by Boreal(15235)

Sarah is x Carmen is x+14
five years ago, they were x-5 and x+9
3(x-5)=x+9
3x-15=x+9
2x=24
x=12
Sarah is 12, Carmen is 26
five years ago, they were 7 and 21 respectively.